# Mobility Edge for the Anderson Model on Random Regular Graphs

### 15 Mar 2026 v1

@@ Appendix A Combes–Thomas Bound
Throughout this appendix, we let $\mathcal{G}=(V,E)$ be a graph (finite or finite) with maximum degree at most $d$, and we retain the notation introduced in Section 6. Additionally, for vertices $v,w$ in a graph, we write $v\sim w$ if there is an edge between $v$ and $w$.
@@ A.1. Finite-Volume Estimate
Fix a finite set $X\subset\mathcal{V}$. For all $r\in\mathbb{N}$ define the (finite) neighborhood
@@ A.2. Passage to Infinite Volume
on $\ell^{2}(\mathcal{V})$ by by cutting all edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ in the underlying graph and removing the corresponding entries from $H$ (so that $\widehat{H}_{r}$ is self-adjoint). If $X,Y\subset\Lambda_{r}$, then
@@
Let $C_{c}(\mathcal{V})$ be the finitely supported functions on $\mathcal{V}$. Since $A$ is bounded and $V$ is a self-adjoint multiplication operator, $C_{c}(\mathcal{V})$ is a core for $H$ and every $\widehat{H}_{r}$. For any $\psi\in C_{c}(V)$, we have $\operatorname{supp}\psi\subset\Lambda_{r}$ for all sufficiently large $r$, in which case the cut edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ do not affect the value of $H\psi$ because $\psi$ vanishes on $\Lambda_{r}^{c}$. Hence for all sufficiently large $r$,
@@
and therefore $\widehat{H}_{r}\psi\to H\psi$ as $r\rightarrow\infty$ for every $\psi\in C_{c}(V)$. By the core criterion for strong resolvent convergence (e.g., [31, Theorem VIII.25]), this implies $\widehat{H}_{r}\to H$ in the strong resolvent sense, which is exactly the desired conclusion. ∎
@@
Let $(T_{r})_{r=1}^{\infty}$ and $T$ be bounded operators on a Hilbert space such that $\sup_{r}\|T_{r}\|<\infty$, and suppose $\lim_{r\rightarrow\infty}T_{r}\to T$ in the strong operator topology. If $P$ and $Q$ are finite-rank orthogonal projections, then
@@
We first prove the claim when both $X$ and $Y$ are finite. Fix finite $X,Y\subset V$ and choose $r$ large enough so that $X\cup Y\subset\Lambda_{r}$. By Lemma A.1 and the identity (A.1),
$\big\|\chi_{X}(\widehat{H}_{r}-z)^{-1}\chi_{Y}\big\|=\big\|\chi_{X}(H_{\Lambda_{r}}-z)^{-1}\chi_{Y}\big\|\leq\frac{2}{\eta}\,e^{-\mu\,\operatorname{dist}(X,Y)}.$ (A.2)
@@
$\|\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}\|\leq C\,e^{-\mu\,\operatorname{dist}(X_{n},Y_{n})}\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}\qquad\text{for all }n.$ (A.3)
@@
Since $(H-z)^{-1}$ is bounded and $\chi_{X_{n}}\to\chi_{X}$, $\chi_{Y_{n}}\to\chi_{Y}$ strongly, both terms on the right-hand side converge to zero in norm. In combination with , this shows that for any $\psi\in\ell^{2}(\mathcal{V})$,
$\|\chi_{X}R(z)\chi_{Y}\psi\|=\lim_{n\to\infty}\|\chi_{X_{n}}R(z)\chi_{Y_{n}}\psi\|\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}\|\psi\|.$
@@
$\|\chi_{X}R(z)\chi_{Y}\|\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}=\frac{2}{\eta}\,e^{-\mu\,\operatorname{dist}(X,Y)}.$
@@ Appendix B Lower Bound on the Density of States
This appendix is devoted to the proof of Lemma 5.6. We begin with an elementary fact about self-adjoint operators.
@@
Let $H$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\varphi\in\mathcal{H}$ such that $\varphi\neq 0$, $E\in\mathbb{R}$, and $\varepsilon>0$. For all Borel sets $J\subset\mathbb{R}$, define.
$\mu_{\varphi}(J)=\langle\varphi,\mathbbm{1}_{J}(H)\varphi\rangle.$
Then
$\mu_{\varphi}([E-\varepsilon,E+\varepsilon])\geq\|\varphi\|^{2}-\frac{\|(H-E)\varphi\|^{2}}{\varepsilon^{2}}.$
@@
The spectral theorem implies
$\|(H-E)\varphi\|^{2}=\int_{\mathbb{R}}(\lambda-E)^{2}\,\mu_{\varphi}(d\lambda)\geq\varepsilon^{2}\,\mu_{\varphi}\big(\mathbb{R}\setminus[E-\varepsilon,E+\varepsilon]\big),$
which yields the conclusion. ∎
For all $M>0$, define the $M$-strength barrier event at distance $1$ from the root of $\mathbb{T}$ by
$\mathcal{B}_{M}=\Big\{\min_{y\sim 0}|V_{y}-V_{0}|\geq M\Big\}.$
On $\mathcal{B}_{M}$, define
$\psi=\delta_{0}+\sum_{y\sim 0}a_{y}\delta_{y},\qquad a_{y}=\frac{t}{V_{y}-V_{0}}.$
Lemma B.2.
On $\mathcal{B}_{M}$,
$\|\psi\|^{2}\leq 1+\frac{dt^{2}}{M^{2}},\qquad\|(H-V_{0})\psi\|^{2}\leq\frac{2t^{4}d^{2}}{M^{2}}.$
@@
By definition, $|a_{y}|\leq t/M$, so
$\|\psi\|^{2}=1+\sum_{y\sim 0}|a_{y}|^{2}\leq 1+\frac{dt^{2}}{M^{2}}.$
Also,
$\big((H-V_{0})\psi\big)(0)=-t\sum_{y\sim 0}a_{y},$
hence $|((H-V_{0})\psi)(0)|\leq t^{2}d/M$.
At any $y\sim 0$, by definition of $a_{y}$ we have
$\big((H-V_{0})\psi\big)(y)=(V_{y}-V_{0})a_{y}-t=0.$
At vertices $u$ with $\operatorname{dist}(u,0)=2$, writing $y$ for the parent of $u$,
$\big((H-V_{0})\psi\big)(u)=-ta_{y},\qquad\big|\big((H-V_{0})\psi\big)(u)\big|\leq\frac{t^{2}}{M}.$
There are exactly d(d1)d(d-1) such vertices, therefore
$\|(H-V_{0})\psi\|^{2}\leq\frac{t^{4}d^{2}}{M^{2}}+\frac{t^{4}d(d-1)}{M^{2}}\leq\frac{2t^{4}d^{2}}{M^{2}}.$
Lemma B.3.
Let $I\subset\mathbb{R}$ be Borel and write $P_{I}=\mathbbm{1}_{I}(H)$. If $\psi=\delta_{0}+\xi$, then
$\mu_{0}(I)\geq\big(\sqrt{\mu_{\psi}(I)}-\|\xi\|\big)^{2}.$
Proof.
Since $\delta_{0}=\psi-\xi$,
$\sqrt{\mu_{0}(I)}=\|P_{I}\delta_{0}\|=\|P_{I}(\psi-\xi)\|\geq\|P_{I}\psi\|-\|P_{I}\xi\|\geq\sqrt{\mu_{\psi}(I)}-\|\xi\|.$
Squaring gives the claim. ∎
Proposition B.4.
Assume $t>0$, fix $M>0$, and set $\varepsilon=4t^{2}d/M$. Then on $\mathcal{B}_{M}$,
$\mu_{0}\big([V_{0}-\varepsilon,V_{0}+\varepsilon]\big)\geq\left(\sqrt{\frac{7}{8}}-\frac{\sqrt{d}\,t}{M}\right)^{2}.$
Proof.
By Lemma B.1 and Lemma B.2,
$\mu_{\psi}([V_{0}-\varepsilon,V_{0}+\varepsilon])\geq\|\psi\|^{2}-\frac{\|(H-V_{0})\psi\|^{2}}{\varepsilon^{2}}\geq 1-\frac{2t^{4}d^{2}/M^{2}}{16t^{4}d^{2}/M^{2}}=\frac{7}{8}.$
Also,
$\xi=\sum_{y\sim 0}a_{y}\delta_{y},\qquad\|\xi\|^{2}=\sum_{y\sim 0}|a_{y}|^{2}\leq\frac{dt^{2}}{M^{2}},$
so $\|\xi\|\leq\sqrt{d}\,t/M$. Apply Lemma B.3. ∎
@@
Fix $E\in[-D,D]$. Set $M=4\sqrt{d}t$ and $\varepsilon=4t^{2}d/M$, and define
$\mathcal{A}_{E,\varepsilon,M}=\big\{|V_{0}-E|\leq\varepsilon/2\big\}\cap\mathcal{B}_{M}.$
On $\mathcal{A}_{E,\varepsilon,M}$ we have $|V_{0}-E|\leq\varepsilon/2$, hence
$V_{0}-\varepsilon,V_{0}+\varepsilon]\subset(E-2\varepsilon,E+2\varepsilon).$
By Proposition B.4, there exists $c_{0}>0$ such that on $\mathcal{B}_{M}$ (hence on $\mathcal{A}_{E,\varepsilon,M}$),
$\mu_{0}\big((E-2\varepsilon,E+2\varepsilon)\big)\geq\mu_{0}\big([V_{0}-\varepsilon,V_{0}+\varepsilon]\big)\geq c_{0}.$
Taking expectation gives
$\nu\big((E-2\varepsilon,E+2\varepsilon)\big)\geq c_{0}\,\mathbb{P}(\mathcal{A}_{E,\varepsilon,M}).$
Observe that
$\mathbb{P}(\mathcal{A}_{E,\varepsilon,M})=\int_{E-\varepsilon/2}^{E+\varepsilon/2}\rho(v)\,\mathbb{P}(\mathcal{B}_{M}\mid V_{0}=v)\,dv.$
Given $V_{0}=v$, the $d$ neighbor potentials are i.i.d. with law $\rho$, so
$\mathbb{P}(\mathcal{B}_{M}\mid V_{0}=v)=\mathbb{P}(|V-v|\geq M)^{d}.$
Lower bounding both factors on $|v-E|\leq\varepsilon/2$ yields
$\mathbb{P}(\mathcal{A}_{E,\varepsilon,M})\geq\varepsilon\Big(\inf_{|v-E|\leq\varepsilon/2}\rho(v)\Big)\Big(\inf_{|v-E|\leq\varepsilon/2}\mathbb{P}(|V-v|\geq M)\Big)^{d}.$
Both factors on the right-hand side can be bounded below by a constant depending only on $d$, $D$, and $\mathfrak{L}$. This completes the proof. ∎

### 19 Mar 2026 v2

@@ Appendix A Combes–Thomas Bound
Throughout this appendix, we let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph (finite or infinite) with maximum degree at most $d$, and we retain the notation introduced in Section 5. Additionally, for vertices $v,w$ in a graph, we write $v\sim w$ if there is an edge between $v$ and $w$.
@@ A.1. Finite-Volume Estimate
Fix any finite set $X\subset\mathcal{V}$ and $r\in\mathbb{N}$, and define the (finite) neighborhood
@@ A.2. Passage to Infinite Volume
on $\ell^{2}(\mathcal{V})$ by cutting all edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ in the underlying graph and removing the corresponding entries from $H$ (so that $\widehat{H}_{r}$ is self-adjoint). If $X,Y\subset\Lambda_{r}$, then
@@
Let $C_{c}(\mathcal{V})$ be the finitely supported functions on $\mathcal{V}$. Since $A$ is bounded and $V$ is a self-adjoint multiplication operator, $C_{c}(\mathcal{V})$ is a core for $H$ and every $\widehat{H}_{r}$. For any $\psi\in C_{c}(\mathcal{V})$, let $\mathcal{C}_{X}\subset\mathcal{V}$ be the union of the connected components of $\mathcal{G}$ that intersect $X$. Since $\operatorname{supp}\psi\cap\mathcal{C}_{X}$ is finite and every vertex in $\mathcal{C}_{X}$ has finite distance to $X$, there exists $R$ such that
$\operatorname{supp}\psi\cap\mathcal{C}_{X}\subset\Lambda_{R-1}.$
Hence for all $r\geq R$,
$\operatorname{supp}\psi\subset\Lambda_{r-1}\cup\mathcal{C}_{X}^{c}.$
Now every edge removed in passing from $H$ to $\widehat{H}_{r}$ has one endpoint in $\Lambda_{r}$ and the other in $\Lambda_{r}^{c}$, so both endpoints lie in $\mathcal{C}_{X}$; moreover, the endpoint in $\Lambda_{r}$ actually lies in $\Lambda_{r}\setminus\Lambda_{r-1}$. Therefore no removed edge is incident to a vertex where $\psi$ is nonzero: on $\mathcal{C}_{X}$ the support of $\psi$ lies inside $\Lambda_{r-1}$, while on $\mathcal{C}_{X}^{c}$ the operators $\widehat{H}_{r}$ and $H$ already agree. Hence for all $r\geq R$,
@@
and therefore $\widehat{H}_{r}\psi\to H\psi$ as $r\rightarrow\infty$ for every $\psi\in C_{c}(\mathcal{V})$. By the core criterion for strong resolvent convergence (e.g., [32, Theorem VIII.25]), this implies $\widehat{H}_{r}\to H$ in the strong resolvent sense, which is exactly the desired conclusion. ∎
@@
Let $(T_{r})_{r=1}^{\infty}$ and $T$ be bounded operators on a Hilbert space such that $\sup_{r}\|T_{r}\|<\infty$, and suppose $T_{r}\to T$ in the strong operator topology as $r\rightarrow\infty$. If $P$ and $Q$ are finite-rank orthogonal projections, then
@@
We abbreviate $\mu=\mu_{\mathrm{CT}}$. First, we prove the claim when both $X$ and $Y$ are finite. Fix finite $X,Y\subset\mathcal{V}$ and choose $r$ large enough so that $X\cup Y_{\mathrm{conn}}\subset\Lambda_{r}$, where $Y_{\mathrm{conn}}$ is the intersection of $Y$ with the union of the connected components of $\mathcal{G}$ meeting $X$. By Lemma A.1 and the identity (A.1),
$\big\|\chi_{X}(\widehat{H}_{r}-z)^{-1}\chi_{Y}\big\|=\big\|\chi_{X}(H_{\Lambda_{r}}-z)^{-1}\chi_{Y_{\mathrm{conn}}}\big\|\leq\frac{2}{\eta}\,e^{-\mu\operatorname{dist}(X,Y)}.$ (A.2)
@@
$\|\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}\|\leq\frac{2}{\eta}\,e^{-\mu\operatorname{dist}(X_{n},Y_{n})}\leq\frac{2}{\eta}\,e^{-\mu\operatorname{dist}(X,Y)}\qquad\text{for all }n.$ (A.3)
@@
Since $(H-z)^{-1}$ is bounded and $\chi_{X_{n}}\to\chi_{X}$, $\chi_{Y_{n}}\to\chi_{Y}$ strongly, both terms on the right-hand side converge to zero in norm. In combination with (A.3), this shows that for any $\psi\in\ell^{2}(\mathcal{V})$,
$\|\chi_{X}(H-z)^{-1}\chi_{Y}\psi\|=\lim_{n\to\infty}\|\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}\psi\|\leq\frac{2}{\eta}e^{-\mu\operatorname{dist}(X,Y)}\|\psi\|.$
@@
$\|\chi_{X}(H-z)^{-1}\chi_{Y}\|\leq\frac{2}{\eta}e^{-\mu\operatorname{dist}(X,Y)}.$
@@ Appendix B Lower Bound on the Density of States
This appendix is devoted to the proof of Lemma 6.2.
For each neighbor $y\sim 0$, let $\mathbb{T}_{y}$ denote the connected component of $\mathbb{T}\setminus\{0\}$ containing $y$, and let $H_{\mathbb{T}_{y}}$ be the restriction of $H_{d,t}$ to $\ell^{2}(\mathbb{T}_{y})$. For $z=E+\mathrm{i}\eta$ with $\eta>0$, define
$\Gamma_{y}(z)=\langle\delta_{y},(H_{\mathbb{T}_{y}}-z)^{-1}\delta_{y}\rangle.$
By the Schur complement formula,
$R_{00}(z)=\langle\delta_{0},(H_{d,t}-z)^{-1}\delta_{0}\rangle=\frac{1}{V_{0}-z-t^{2}\sum_{y\sim 0}\Gamma_{y}(z)}.$
@@
There exists a constant $p_{*}>0$ such that for all $E\in\mathbb{R}$, all $\eta>0$, and all neighbors $y\sim 0$,
$\mathbb{P}\big(|\Gamma_{y}(E+\mathrm{i}\eta)|\leq 1\big)\geq p_{*}.$
@@
Fix $z=E+\mathrm{i}\eta$ and $y\sim 0$. Conditioning on the potentials in $\mathbb{T}_{y}\setminus\{y\}$, the Schur complement formula gives
$\Gamma_{y}(z)=\frac{1}{V_{y}-a-\mathrm{i}b}$
for some random variables $a\in\mathbb{R}$ and $b>0$ that are independent of $V_{y}$. Hence
$\mathbb{P}\big(|\Gamma_{y}(z)|\leq 1\,\big|\,a,b\big)=\mathbb{P}\big((V_{y}-a)^{2}+b^{2}\geq 1\,\big|\,a,b\big)\geq\mathbb{P}\big(|V_{y}-a|\geq 1\big).$
Since the density of $V_{0}$ is even and strictly decreasing on $0,\infty)$, among all intervals of length $2$ the interval [1,1][-1,1] has the largest Gaussian mass. Therefore
$\mathbb{P}(|V_{y}-a|\geq 1)\geq\mathbb{P}(|V_{y}|\geq 1)>0$
for all $a\in\mathbb{R}$. Setting $p_{*}=\mathbb{P}(|V_{y}|\geq 1)$ and averaging over $a,b$ proves the claim. ∎
Proposition B.2.
Fix $d\geq 3$, $t>0$, and $D>0$. Then there exists a constant $c_{0}(d,t,D)>0$ such that for all $E\in[-D,D]$ and all $\eta\in(0,1]$,
$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(E+\mathrm{i}\eta)\geq c_{0}.$
@@
Fix $E\in[-D,D]$ and $\eta\in(0,1]$, and set $z=E+\mathrm{i}\eta$. Define the event
$\mathcal{G}_{z}=\bigcap_{y\sim 0}\big\{|\Gamma_{y}(z)|\leq 1\big\}.$
Since the forward subtrees $(\mathbb{T}_{y})_{y\sim 0}$ are disjoint and the potentials are i.i.d., the random variables $(\Gamma_{y}(z))_{y\sim 0}$ are independent. Hence, by Lemma B.1,
$\mathbb{P}(\mathcal{G}_{z})\geq p_{*}^{d}.$
Conditioning on the sigma-field generated by $(V_{x})_{x\neq 0}$, define
$A_{z}=E+t^{2}\sum_{y\sim 0}\operatorname{Re}\Gamma_{y}(z),\qquad B_{z}=\eta+t^{2}\sum_{y\sim 0}\operatorname{Im}\Gamma_{y}(z).$
Then $B_{z}>0$ and
$R_{00}(z)=\frac{1}{V_{0}-A_{z}-\mathrm{i}B_{z}}.$
Therefore,
$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(z)=\frac{1}{\pi}\mathbb{E}\big[\operatorname{Im}R_{00}(z)\big]=\mathbb{E}\!\left[\int_{\mathbb{R}}\rho(v)\,\frac{1}{\pi}\frac{B_{z}}{(v-A_{z})^{2}+B_{z}^{2}}\,dv\right],$
where $\rho$ is the density of $V_{0}$. On the event $\mathcal{G}_{z}$ we have
$|A_{z}|\leq D+t^{2}d,\qquad 0<B_{z}\leq 1+t^{2}d.$
Set
$K_{1}=D+t^{2}d,\qquad K_{2}=1+t^{2}d.$
Define, for $a\in\mathbb{R}$ and $b>0$,
$F(a,b)=\int_{\mathbb{R}}\rho(v)\,\frac{1}{\pi}\frac{b}{(v-a)^{2}+b^{2}}\,dv.$
Because $\rho$ is continuous and strictly positive, $F$ extends continuously to $\mathbb{R}\times[0,\infty)$ by setting $F(a,0)=\rho(a)$ (see, e.g., [13, Theorem 7.4]). Hence
$\inf_{|a|\leq K_{1},\;0\leq b\leq K_{2}}F(a,b)>0.$
Denote this infimum by $c_{1}$. On $\mathcal{G}_{z}$ we have $|A_{z}|\leq K_{1}$ and $0<B_{z}\leq K_{2}$, so
$F(A_{z},B_{z})\geq c_{1}.$
Consequently,
$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(z)=\mathbb{E}\big[F(A_{z},B_{z})\big]\geq c_{1}\,\mathbb{P}(\mathcal{G}_{z})\geq c_{1}p_{*}^{d}.$
Thus the claim holds with $c_{0}=c_{1}p_{*}^{d}$. ∎
@@
By Lemma 6.1, the measure $\nu$ is absolutely continuous with respect to Lebesgue measure, so there exists $n\in L^{\infty}(\mathbb{R})$ such that
$\nu(d\lambda)=n(\lambda)\,d\lambda.$
For $\eta>0$, let
$P_{\eta}(x)=\frac{1}{\pi}\frac{\eta}{x^{2}+\eta^{2}}$
denote the Poisson kernel. Since $m_{\mathbb{T}}$ is the Stieltjes transform of $\nu$, we have
$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(E+\mathrm{i}\eta)=\int_{\mathbb{R}}P_{\eta}(E-\lambda)\,n(\lambda)\,d\lambda=(P_{\eta}*n)(E).$
By Proposition B.2, for every $E\in[-D,D]$ and every $\eta\in(0,1]$,
$(P_{\eta}*n)(E)\geq c_{0}.$
Now fix an interval $I\subset[-D,D]$. Integrating over $E\in I$ gives
$\int_{I}(P_{\eta}*n)(E)\,dE\geq c_{0}|I|\qquad\text{for all }\eta\in(0,1].$
Since $(P_{\eta})_{\eta>0}$ is an approximate identity and $n\in L^{\infty}(\mathbb{R})$, we have
$\lim_{\eta\downarrow 0}\int_{I}(P_{\eta}*n)(E)\,dE=\int_{I}n(E)\,dE=\nu(I).$
Therefore, $\nu(I)\geq c_{0}|I|$, as desired. ∎